Tags:
create new tag
view all tags

Shift plan 14/12/2011

Access: 1 - 1.5 hours

  • Low energy green alignment post IP - remove new recollimating lens, replace turning mirror on stage, turn beam onto ceramic beam dump.
  • Check reasonable laser signal on APD.
  • High energy laser transmission check through IP (measure with energy meter + attenuator).
  • Reference laser position with screen (including horizontal movement check).
  • Replace final focus lens with undamaged one - check if threads on lens mount need lubricating.
  • Get red alignment laser down beampipe to check OTR alignment.

Beam time: 2 hours max.

  • Check background level (previous week ~ 1V) and tune for low background, aiming for 500mV.

Beam time: 30 - 45 minutes

  • OTR images with < 700ms exposure time, calibration, quick quad scan to minimise beam size.

Beam time: remaining ~ 4 hours

  • Laserwire scans - timing, alignment, vertical scans (~ 2 um steps), then if no signal alter beam pointing with magnets to try and hit window.
  • If signal, quad scans, telescope scans, lots of data taking.

Physics WebpagesRHUL WebpagesCampus Connect • Royal Holloway, University of London, Egham, Surrey TW20 0EX; Tel/Fax +44 (0)1784 434455/437520

Topic revision: r1 - 13 Dec 2011 - LauraCorner

 
This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding RHUL Physics Department TWiki? Send feedback